# GLL: A Differentiable Graph Learning Layer for Neural Networks

### v1

@@ 4.3.3 Adversarial Experiments
Figure 7: Adversarial robustness to FGSM, IFGSM, and CW attacks on the CIFAR-10 dataset. The backbone network is the same in all cases (ResNet-18); the only difference is the classifier (either MLP with softmax or our novel GLL). We report results for both robustly and naturally trained models.
We further investigate the performance of GLL against CW attacks in Figure 8. This figure plots the accuracy versus the average distance (squared) that the adversary moves an image in the test set (instead of the value of $c$). This reveals a striking difference between MLP and GLL: GLL resists the CW adversary from moving the images significantly more than the MLP classifier, making it more robust to these attacks. The one exception is the FashionMNIST robust GLL model, which still outperforms the MLP-based model across distances and values of $c$.
@@
Figure 8: Comparison of accuracy vs. average distance of attack for CW attack on both datasets. Each point represents a different value of $c$. We see that GLL resists the adversary from moving the target images significantly more than the standard MLP classifier.
We see that, across datasets, training methods (both natural and robust), severity of attacks, and neural network architectures, our graph learning layer is significantly more robust to adversarial attacks, notably without sacrificing in natural accuracy. Laplace learning can be viewed as repeatedly computing a weighted average over a given node’s neighbors, and this relational information may be preventing attacks from successfully perturbing the input images to fool the network. In particular, we note that the models employing the standard classifier quickly decay to 0 accuracy against IFGSM and CW attacks, while GLL shows robustness to these attacks, even for extremely large values of $\epsilon$ and $c$, respectively. Moreover, performance of our GLL could likely be even further improved with more computational resources, as graph learning techniques tend to improve as the number of nodes (i.e. the batch size) increases.
@@ 5 Conclusion
Graph-based learning is a powerful class of machine learning techniques that have been developed throughout the 21st century. As research in deep learning has exploded throughout the same time period, there has been significant interest in integrating deep learning with similarity matrix-based graph learning. This paper gives a precise mathematical framework to do exactly that. We derived the necessary gradient backpropagation equations for general nonlinear elliptic graph Laplace equations, allowing us to fully integrate graph-based classifiers into the training of neural networks. Using this novel framework, we introduced the Graph Learning Layer, or GLL, which replaces the standard projection head and softmax classifier in classification tasks. Experimentally, we demonstrated how GLL improves upon the standard classifier in several ways: it can learn the geometry underlying data, it can improve the training dynamics and test accuracy on benchmark datasets in both small and large networks, and it is significantly more robust to a variety adversarial attacks in both naturally and robustly trained models.
There are many exciting avenues of future work in this area. Given the strong empirical results, theoretically guaranteeing the adversarial robustness of GLL is an important extension of this work. Graphs are a common setting for active learning techniques Settles (2009); Miller and Bertozzi (2024); the full integration of graph learning into a neural network provides a natural way to do active learning in a deep learning context. Our theoretical work also provides a way to integrate other graph learning algorithms - such as $p$-Laplace learning and Poisson learning - into a deep learning framework, and numerical experiments could yield similarly strong results as reported in this paper. Many performant semisupervised deep learning techniques utilize psuedolabels generated by graph learning methods Iscen et al. (2019); Sellars et al. (2021); we hope the methods presented here will lead to further improvements and insights.
@@ Appendix A Training Details
Prior to GLL-based fully supervised training for the CIFAR-10 dataset, we initialize the network’s feature encoder using pre-trained weights from SimCLR (Chen et al. (2020a, b)). SimCLR is a framework for contrastive learning of visual representations. For a neural network $f$ and a minibatch of $m$ samples $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{m}\}$, each $\mathbf{x}_{k}$ is augmented into pairs $\tilde{\mathbf{x}}_{2k-1},\tilde{\mathbf{x}}_{2k}$. These are processed through the network: $\mathbf{z}_{2k-1}=f(\tilde{\mathbf{x}}_{2k-1}),\mathbf{z}_{2k}=f(\tilde{\mathbf{x}}_{2k})$. The SimCLR loss is defined as:
$\begin{split}&\mathcal{L}=\frac{1}{2m}\sum_{k=1}^{m}[\ell(2k,2k-1)+\ell(2k-1,2k)],\\ &\ell(i,j)=-\log\frac{\exp(g(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1,k\neq i}^{2m}\exp(g(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)},\end{split}$ (46)
where $\tau$ is a temperature parameter, and $g(\mathbf{z}_{i},\mathbf{z}_{j})=\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/(\|\mathbf{z}_{i}\|\|\mathbf{z}_{j}\|)$ is the cosine similarity. It is worth noting that the cosine similarity used in the SimCLR loss (46) closely resembles the similarity measure we employ when constructing graphs in graph Laplace learning. This similarity suggests that the feature vectors generated by a network trained with SimCLR already possess a favorable clustering structure. Consequently, utilizing SimCLR to pretrain the feature encoder in our network can significantly reduce the difficulty of subsequent training stages.
In the GLL-based fully supervised training process, our approach differs from classic training methods. For each minibatch, in addition to batch images and their corresponding labels, we also require base images and their associated base labels. While this base data can remain fixed throughout the entire training process, updating the base dataset during training can enhance the overall effectiveness. The base dataset in graph Laplace learning represents the labeled nodes, which can also be viewed as the sources in a label propagation process. An ideal base dataset should be relatively dispersed and distributed near the decision boundaries of different clusters.
Let $X$ denote the training dataset, and $B$ be the number of base data points. We can simply update the based dataset as a random subset $X_{b}\subset X$ such that $|X_{b}|=B$. Or we can define a score to evaluate which data is more beneficial if selected. Assume that at a certain epoch, the Laplace learning output for $\mathbf{x}_{j}\in X\setminus X_{b}$ is $\mathbf{u}_{j}=(u_{j}^{1},u_{j}^{2},\ldots,u_{j}^{C})$, where $C$ is the total number of classes. We can define two types of scores that essentially measure the uncertainty of the output $\mathbf{u}_{j}$:
@@
We then select the base dataset $X_{b}$ as the data with the highest scores. However, there’s a possibility that the selected high-score data points may cluster together in the graph structure, potentially providing redundant information. A better approach is to employ the LocalMax batch active learning method (Chapman et al. (2023); Chen et al. (2024)), which selects nodes in the graph that have high scores and satisfy a local maximum condition. By doing so, we can avoid selecting nodes too close to each other, ensuring a more diverse and informative base dataset.
@@ B.3 Further CW Experiments
Figure 11: Comparison of accuracy vs. average distance of attack for CW attack on both datasets. Each point represents a different value of $c$, where we plot additional values of $c$ for the GLL-based models compared to Figure 8. We see that GLL is still superior, even when controlling for distance.

### v2

@@ 4.6.3 Adversarial Experiments
Figure 8: Adversarial robustness to FGSM, IFGSM, and CW attacks on the CIFAR-10 dataset. The backbone network is the same in all cases (PreActResNet-18); the only difference is the classifier (either MLP with softmax, WNLL, or our novel GLL). We report results for both robustly and naturally trained models.
Taken together, our results (Table 10 and Figures 6, 7, and 8) demonstrate that across datasets, training methods (both natural and robust), severity of attacks, and neural network architectures, models trained with GLL are significantly more robust to adversarial attacks, without sacrificing natural accuracy. Laplace learning can be viewed as repeatedly computing a weighted average over a given node’s neighbors, and this relational information may be preventing attacks from successfully perturbing the input images to fool the network compared to the MLP classifier models. Moreover, our results suggest that tracking the gradients exactly (GLL), rather than approximating them (WNLL), offers a significant increase in robustness. GLL’s advantage is even more apparent for stronger attacks, whereas MLP and WNLL-based models decay much more quickly toward random guessing or even 0% accuracy. Moreover, performance of our GLL could likely be even further improved with more computational resources, as graph learning techniques tend to improve as the number of nodes (i.e. the batch size) increases.
@@
Figure 9: Comparison of accuracy vs. average distance of attack for CW attack on all three datasets. Each point represents a different value of $c$. We see that GLL resists the adversary from moving the target images significantly more than the standard MLP or WNLL classifier, suggesting that graph-based classification and precise tracking of gradients is key for adversarial robustness.
We further investigate the improved robustness of GLL against CW attacks in Figure 9 (see also Appendix B.3). This figure plots the accuracy versus the average distance (squared) that the adversary moves an image in the test set (instead of the value of $c$). This reveals a striking difference: models trained with the GLL classifier resist the CW adversary from moving the images significantly more than the MLP or WNLL-based models, resulting in more robustness to these attacks. Interestingly, WNLL is the least resistant to these perturbations. Since WNLL uses linear layer gradients and GLL uses precise gradients, these results suggest that the exact backpropagation in GLL aids robustness to these attacks. One hypothesis is that the relational information in GLL’s gradients “trap” data points in neighborhoods with points of the same class. Theoretically guaranteeing GLL’s adversarial robustness is an interesting line of future work.
@@ 5 Conclusion
Graph-based learning is a powerful class of machine learning techniques that have been developed throughout the 21st century. As research in deep learning has exploded throughout the same time period, there has been significant interest in integrating deep learning with similarity matrix-based graph learning. This paper gives a precise mathematical framework to do exactly that. We derived the necessary gradient backpropagation equations for general nonlinear elliptic graph Laplace equations, allowing us to fully integrate graph-based classifiers into the training of neural networks. Using this novel framework, we introduced the Graph Learning Layer, or GLL, which replaces the standard projection head and softmax classifier in classification tasks. Experimentally, we demonstrated how the GLL improves upon the standard classifier and previous graph-based layers (WNLL) in several ways. Two illustrative examples demonstrated that GLL-based networks can learn the geometry of the underlying data and improve training dynamics in overparameterized settings, respectively. In a suite of large scale experiments, we showed that networks trained with a GLL classifier have improved generalization on benchmark datasets, especially when training labels are limited. Although the improvement in natural accuracy was occasionally modest, later experiments demonstrated the GLL’s significant and consistent performance gains for networks in the adversarial setting. This points toward a practical advantage of GLL in safety- and security-critical settings. Moreover, by comparing to WNLL-based networks, our results also empirically validated the advantage of tracking gradients through graph learning precisely, rather than approximately.
There are many exciting avenues of future work in this area. Given the strong empirical results, theoretically guaranteeing the adversarial robustness of the GLL is an important extension of this work. Graphs are a common setting for active learning techniques (settles2009active; miller2024model); the full integration of graph learning into a neural network provides a natural way to do active learning in a deep learning context. Our theoretical work also provides a way to integrate other graph learning algorithms - such as $p$-Laplace learning and Poisson learning - into a deep learning framework, and numerical experiments could yield similarly strong results as reported in this paper. Many performant semi-supervised deep learning techniques utilize psuedolabels generated by graph learning methods (iscen2019label; sellars2021laplacenet); we hope the methods presented here will lead to further improvements and insights.
@@ Appendix A Training Details
A.1 Sampling from Base and Unlabeled Datasets
In the GLL-based training process, our approach to constructing mini-batches differs from classical fully supervised training. As described in Section 4.1.1, each mini-batch $\mathcal{B}$ contains a base dataset $\mathcal{L}_{b}\subset\mathcal{L}$ together with additional labeled and unlabeled samples $\mathcal{B}_{l}$ and $\mathcal{B}_{u}$. For each mini-batch, in addition to the images whose labels are used to compute the loss (the samples in $\mathcal{B}_{l}$), we also require base images and their associated base labels (the samples in $\mathcal{L}_{b}$), which act as labeled nodes for graph-based label propagation. While this base dataset can be kept fixed throughout training, we also consider two update strategies for $\mathcal{L}_{b}$ that can further enhance performance.
First, in the random resampling strategy, we update the base dataset at the beginning of each epoch by sampling a random subset $\mathcal{L}_{b}\subset\mathcal{L}$ with $|\mathcal{L}_{b}|=N_{b}$. To provide stable supervision for label propagation, $\mathcal{L}_{b}$ is chosen so that the class proportions in $\mathcal{L}_{b}$ match the class proportions in the full labeled set $\mathcal{L}$, for example by stratified sampling.
Second, in an uncertainty-based strategy, we explicitly score labeled samples by how informative they would be if selected into $\mathcal{L}_{b}$. In the graph Laplace learning framework, the base dataset $\mathcal{L}_{b}$ represents the labeled nodes and can be viewed as the sources in a label propagation process. An ideal base dataset should be relatively dispersed in the graph and lie near decision boundaries between clusters, so that it provides informative supervision to the unlabeled nodes.
Let $\mathcal{X}=\mathcal{L}\cup\mathcal{U}$ denote the full training set, where $\mathcal{L}$ is the labeled subset and $\mathcal{U}$ is the unlabeled subset, and let $N_{b}$ be the desired size of the base dataset. Assume that at a given epoch, the GLL prediction for a labeled sample $\mathbf{x}_{j}\in\mathcal{L}\setminus\mathcal{L}_{b}$ is
$\mathbf{u}_{j}=\bigl(u_{j}^{1},u_{j}^{2},\ldots,u_{j}^{C}\bigr),$
where $C$ is the number of classes and $u_{j}^{c}$ is the predicted confidence for class $c$. We can define two uncertainty-based scores for $\mathbf{u}_{j}$:
@@ A.1 Sampling from Base and Unlabeled Datasets
Both scores are larger when the prediction $\mathbf{u}_{j}$ is more uncertain. In this second strategy, we select the new base dataset $\mathcal{L}_{b}$ as the $N_{b}$ labeled samples in $\mathcal{L}$ with the highest scores, according to either (57) or (58).
@@
However, there is a risk that the highest-scoring samples may cluster in a small region of the graph, providing redundant information. A more effective variant is to employ the LocalMax batch active learning method (chapman2023novel; chen2024batch). In this approach, nodes in the graph are first assigned uncertainty scores, and then a subset is selected that both has high scores and satisfies a local maximum condition in the graph. This discourages the selection of nodes that are too close to each other and leads to a more diverse and informative base dataset $\mathcal{L}_{b}$ for GLL.
A.2 Contrastive Learning and Warm Up
Prior to GLL-based fully supervised training for the CIFAR-10 dataset, we initialize the network’s feature encoder using pre-trained weights from SimCLR (chen2020simple; chen2020big). SimCLR is a framework for self-supervised contrastive learning of visual representations. For a neural network $f$ and a minibatch of $m$ samples $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{m}\}$, each $\mathbf{x}_{k}$ is augmented into pairs $\tilde{\mathbf{x}}_{2k-1},\tilde{\mathbf{x}}_{2k}$. These are processed through the network: $\mathbf{z}_{2k-1}=f(\tilde{\mathbf{x}}_{2k-1}),\mathbf{z}_{2k}=f(\tilde{\mathbf{x}}_{2k})$. The SimCLR loss is defined as:
$\begin{split}&\mathcal{L}_{\mathrm{sim}}=\frac{1}{2m}\sum_{k=1}^{m}[\ell(2k,2k-1)+\ell(2k-1,2k)],\\ &\ell(i,j)=-\log\frac{\exp(g(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1,k\neq i}^{2m}\exp(g(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)},\end{split}$ (59)
where $\tau$ is a temperature parameter, and $g(\mathbf{z}_{i},\mathbf{z}_{j})=\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/(\|\mathbf{z}_{i}\|\|\mathbf{z}_{j}\|)$ is the cosine similarity. Essentially, SimCLR teaches the network to increase the similarity between embedded pairs and reduce the similarity between the other embedded pairs. It is worth noting that the cosine similarity used in the SimCLR loss (59) closely resembles the similarity measure we employ when constructing graphs in graph Laplace learning. This similarity suggests that the feature vectors generated by a network trained with SimCLR already possess a favorable clustering structure. Consequently, utilizing SimCLR to pretrain the feature encoder in our network can significantly reduce the difficulty of subsequent training stages.
In the supervised setting, SupCon (khosla2020supervised) can be used instead of SimCLR. With full access to the underlying labels, the goal is now to maximize the similarity between all embedded pairs within the same class and minimize the similarity of samples across classes. The loss function is
$\mathcal{L}_{\mathrm{sup}}=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp\left(z_{i}\cdot z_{p}/\tau\right)}{\sum_{a\in A(i)}\exp\left(z_{i}\cdot z_{a}/\tau\right)}$ (60)
where $P(i)\equiv\left\{p\in A(i):\tilde{\boldsymbol{y}}_{p}=\tilde{\boldsymbol{y}}_{i}\right\}$ is the set of indices for all samples that belong to the same class as sample $i$ (excluding itself) and $A(i)$ is the set containing all indices except $i$.
For the purposes of warming up the GLL, either SimCLR or SupCon can be used depending on the framework. In our work, we have even seen success in balancing a combination of the two methods to utilize the labels for a strong embedding while not overfitting too strongly. The pretraining loss is taken to be the same mixed objective as in (53),
$\mathcal{L}_{\mathrm{pretrain}}=\gamma\mathcal{L}_{\mathrm{sim}}+(1-\gamma)\mathcal{L}_{\mathrm{sup}}$
for some $\gamma\in[0,1]$.
A.3 Data Augmentation Details
We describe here the data augmentation policies used during both contrastive pretraining and supervised training.
Augmentation pool for color images.
For color image datasets, we adopt a policy similarly to the RandAugment cubuk2019randaugment. Let $\mathcal{T}$ denote a pool of primitive transformations including
  • Photometric operations: autocontrast, histogram equalization, brightness, contrast, color, sharpness, posterization, solarization;
  • Geometric operations: rotations, horizontal shears and vertical shears, horizontal translations and vertical translations;
  • The identity transform.
For each image in a mini-batch, we randomly sample a fixed number of transformations from $\mathcal{T}$ with replacement and apply them sequentially. Each selected transform is assigned a magnitude drawn uniformly from a transform-specific range. For example, rotation angles up to a few tens of degrees, shear factors bounded by a small absolute value, or contrast/brightness factors within a bounded interval. After these operations, we apply a cutout transform that removes a randomly located square patch of the image, with the side length sampled as a fraction of the image size. All sampling is done independently for each image and for each view, so that different augmented views of the same sample are stochastically decorrelated.
Augmentation pool for grayscale images.
For grayscale datasets, we use a variant of the above policy tailored to single-channel images. In this case the transformation pool contains only operations that are meaningful on grayscale inputs:
  • Geometric transforms: rotations, shears, and translations in the horizontal and vertical directions;
  • Photometric transforms: inversion, histogram equalization, solarization, brightness, contrast, and sharpness adjustments;
  • Cutout with a square mask whose side length is sampled up to a fixed proportion, filled with a constant gray or black intensity.
As in the color case, we randomly select a small number of transforms for each image, sample their magnitudes from predefined ranges, and apply them sequentially. The parameters of the grayscale policy such as the number of operations, the global magnitude index, and optional magnitude jitter are shared across the dataset, while the specific operations and their sampled magnitudes vary independently from image to image.
Usage in pretraining and supervised training.
During contrastive pretraining, the above policies are used to create multiple stochastic views of each input example. All samples in a pretraining mini-batch participate in the SimCLR loss via their augmented views, while only the labeled samples in the batch participate in the SupCon loss. In the subsequent supervised training stage with GLL, we continue to apply the same augmentation pipeline to every training sample before feeding it through the encoder and the classifier. In this way, both the contrastive pretraining and the supervised training phases benefit from the same set of strong yet label-preserving data transformations.
@@ B.3 Further CW Experiments
Figure 12: Comparison of accuracy vs. average distance of attack for CW attack on both datasets. Each point represents a different value of $c$, where we plot additional values of $c$ for the GLL-based models compared to Figure 9. We see that GLL is still superior at resisting perturbations by CW attack, even when $c$ is orders of magnitude larger than for WNLL or MLP, and even when controlling for distance on CIFAR-10.